# Q-SLAM: Quadric Representations for Monocular SLAM

### v1

@@ Abstract
Monocular SLAM has long grappled with the challenge of accurately modeling 3D geometries. Recent advances in Neural Radiance Fields (NeRF)-based monocular SLAM have shown promise, yet these methods typically focus on novel view synthesis rather than precise 3D geometry modeling. This focus results in a significant disconnect between NeRF applications, i.e., novel-view synthesis and the requirements of SLAM. We identify that the gap resulting from the volumetric representations used in NeRF, which are often dense and noisy. In this study, we propose a novel approach that reimagines volumetric representations through the lens of quadric forms. We posit that most indoor scene components can be effectively represented as quadric planes. Leveraging this assumption, we reshape the volumetric representations with million of cubes by several quadric planes, which leads to more accurate and efficient modeling of 3D scenes in SLAM contexts. Our method involves two key steps: First, we use the quadric assumption to enhance noisy depth estimations obtained from the tracking module. This step alone significantly improves depth estimation accuracy. Second, in the subsequent mapping phase, we diverge from previous NeRF-based SLAM methods that distribute sampling points across the entire volume space. Instead, we concentrate sampling points around quadric planes and aggregate them using a novel quadric-decomposed Transformer. The quadric semantics are not only explicitly used for depth correction and scene decomposition, but also serve as an implicit supervision signal during the training of NeRF network. Additionally, we introduce an end-to-end joint optimization strategy that synchronizes pose estimation with 3D reconstruction. Through rigorous experimental evaluation, our method exhibits superior performance over other approaches relying on estimated depth. Furthermore, our method achieves accuracy comparable to methods utilizing ground truth depth, marking a significant advancement in the field of monocular SLAM and the integration of NeRF-based techniques for accurate 3D geometry modeling and joint optimization of camera poses.
Keywords:
Neural Radiance Fields Simultaneous Localization and Mapping
@@ 1 Introduction
(a) Volumetric representation
(b) Quadric representation
Figure 1: Comparison between the volume representation and our quadric representation. With an off-the-shelf segmentation network, we can partition the 3D scenes into different quadric surfaces. Instead of processing every points independently, we can fully exploit the correlation between points on and across the quadric surfaces.
@@
To enable high-quality 3D scene reconstruction, recent SLAM approaches resort to NeRFs and Gaussians as key map representations. Specifically, NeRFs [14], employing neural networks i.e. MLPs, are recognized for their ability to achieve photo-realistic novel-view synthesis, as demonstrated in studies such as iMAP [30], Nice-SLAM [52]. Conversely, 3D Gaussian Splatting (3DGS) [11] relies on dense Gaussian representations to learn scene geometries effectively. Despite achieving photo-realistic novel-view synthesis, NeRFs and Gaussians are not necessarily compatible in SLAM systems, with several drawbacks. On the one hand, MLP-based NeRFs are slow to train while achieving significant memory efficiency. On the other hand, 3D Gaussian Splatting offers near-real-time training capabilities but suffer from large memory consumption and sensitivity to inaccurate camera pose estimation. Additionally, both methods focus on dense high-quality reconstruction, demanding substantial representation efforts to capture fine-grained details.
@@
Addressing these challenges necessitates the design of map representations that are compatible with both learning-based localization front-end and NeRF(GS)-based reconstruction module. Therefore, to enhance scene geometry representation in SLAM, we introduce a novel method, termed Q-SLAM. This approach begins with a fundamental premise: most scene elements, particularly in indoor environments, are effectively representable by quadric planes, as illustrated in Figure 1. Leveraging this assumption, we integrate the proposed quadric representations across the SLAM pipeline in several key areas.
  1. 1.
@@
    The well-established front-end tracking module can take RGB stream as input and offer a rough depth estimation while estimating camera poses. Nevertheless, this depth estimation encounters significant challenges, particularly at edges and in texture-less regions, where accuracy tends to degrade. To address these issues, we introduce a quadric- based depth correction module, leveraging quadric surfaces to refine the depth estimation, thus improving overall scene reconstruction accuracy, especially in complex and challenging environments.
  2. 2.
@@
    In the mapping module, departing from previous NeRF-based SLAM approaches like [22, 52], which rely on volumetric representations, our method transforms densely volumetric scenes - comprising millions of cubes (e.g., $300\times 300\times 300$ voxels) - into a more manageable set of quadric surfaces (50-100 quadric surfaces per keyframe). This approach maintains the integrity of the 3D scene descriptions while significantly reducing complexity, by sampling around the quadric surfaces instead of the whole 3D space.
  3. 3.
@@
    During the rendering phase, we develop a quadric-ray transformer. We employ importance-sampling based on the quadric-rectified depth values, and we emphasize that sampled points on a quadric naturally belong to the same instance, which facilitates the modeling of their interrelationships during rendering, akin to the method in [38, 12]. We first perform feature interaction between points along a ray to aggregate information and then model the relationships between these surface points across rays using a transformer.
  4. 4.
@@
    Notably, we extend the utility of quadric semantics beyond mere depth correction and scene decomposition. They are also used as a supervision signal throughout the optimization of the NeRF network to learn robust and accurate features. With the integration of quadrics into the pipeline, we also consider the estimated pose as a learnable parameter in our mapping phase. This introduces a differentiable optimization process to refine the pose estimation, enhancing overall system accuracy and efficiency.
We conduct experiments on Replica [28], ScanNet [6], and TUM RGB-D [29], and evaluate the results on novel-view synthesis, depth estimation, and pose estimation tasks. The results demonstrate the effectiveness of Q-SLAM, with 6.9 % PSNR (for novel-view synthesis) and 13.2 % L1 error (for depth estimation) improvements on Replica dataset, 4.1 %, 5.7 % ATE (for pose estimation) improvements on ScanNet, and TUM RGB-D datasets, respectively.
@@ 2.1 Dense Visual SLAM
Neural Radiance Fields (NeRF) [13] have exhibited notable efficacy in diverse applications, including view synthesis [13, 38, 37] and 3D reconstruction [4, 2]. Concurrently, Simultaneous Localization and Mapping (SLAM) confronts the challenge of precisely determining the pose of a moving sensor within an environment. The output pose of SLAM aligns seamlessly with the input requirements of NeRF, making NeRF a natural choice as the mapping backbone in a SLAM system. RGBD SLAM: Innovative contributions such as iMAP [30] and Nice-SLAM [52] have pioneered the integration of NeRF into SLAM systems, demonstrating commendable tracking and mapping performance for indoor scenes. Co-SLAM [36] adopts a multi-resolution hash-grid representation for the scene to expedite convergence. E-SLAM [9] integrates the multi-scale axis-aligned perpendicular feature planes for efficient reconstruction of 3D scenes. By anchoring neural scene features in a point cloud generated iteratively in an input-driven manner, Point-SLAM [25] optimizes runtime and memory usage while maintaining fine detail resolution. Besides using NeRF as the map representation, recent SLAM systems also resort to 3D Gaussian Splatting (3DGS) [11] due to its explicit representation property and real-time rendering performance. SplaTAM [10] pioneers in replacing NeRF with 3DGS and has achieved promising results in both tracking and mapping. GS-SLAM [43] further proposes a coarse-to-fine technique to select reliable 3D Gaussian representations for camera pose optimization effectively. However, these GS-based approaches have one drawbacks in common, much higher requirements for memory usage.
All these approaches above rely on ground truth depth map for supervision of NeRF (or 3DGS) training, which is hard to obtain in real-world environment, thereby limiting their applicability in robotics. Conversely, Monocular SLAM method emerges, with only RGB images as input. In particular, NeRF-SLAM [22] integrates DROID-SLAM [34] for pose estimation and depth prediction and use Instant-NGP [16] to fit a NeRF. GO-SLAM [49] globally optimize poses and 3D reconstruction. However, the performance of these methods in 3D reconstruction is impeded by the inherent limitations of noisy predicted depth maps.
@@ 2.2 Scene representations
Volumetric Representations are predominantly utilized in Neural Radiance Fields (NeRFs) [3, 17, 13], and in the 3D perception field [23, 41, 21]. These representations encapsulate a 3D scene through a volumetric approach, where each volume element embodies occupancy probabilities or task-specific features. However, deriving volumetric representations solely from visual inputs presents significant challenges, particularly in terms of geometric estimation, leading to ambiguous and noisy outputs. Point-Cloud Representation offers a sparser alternative [42]. Point-SLAM [25] utilize a dynamic point density strategy to reduce computational and memory usage. Yet sparse point-clouds struggle to depict scenes comprehensively, while dense point-clouds encounter similar complications to volumetric representations, as previously mentioned. Consequently, few vision-based SLAM systems adopt point-based representations. Additionally, primitives like Planes [7] and Quadric-Meshes [15, 8, 31] have been proposed for scene representation. QuadricSLAM [20] estimates 3D quadric surfaces from 2D multi-view images and uses them as landmark representations for camera tracking in the SLAM pipeline. Dense Planar SLAM [24] uses bounded planes and surfels extracted from depth images to build a real-time SLAM system, but ignore the fine-grained reconstruction of 3D scenes. Point-plane SLAM [48] exploits the constraints of plane edges, which are predominant features that are less affected by measurement noise. ManhattanSLAM [45] utilizes the indoor structural information to estimate camera poses based on the Manhattan World (MW) assumption. However, these are typically employed for regularization during optimization or are challenging to optimize due to the need to train parameters from scratch, limiting their applicability in visual SLAM.
Drawing inspiration from quadric representations in LiDAR SLAM [40], we introduce quadric representations in visual monocular SLAM. Instead of merely using quadric primitives for regularization, we integrate them comprehensively into the entire SLAM pipeline, including depth correction, the innovative quadric ray transformer and the quadric semantics as supervision signal.
@@ 3.1 Overview
Our framework, Q-SLAM, as illustrated in Fig. 2, takes monocular RGB sequences as input. Initially, the tracking module predicts rough depth maps and initial camera poses from these inputs. Concurrently, a pretrained segmentation network is employed to estimate segmentation masks from these images. These masks are then utilized by the quadric-based depth correction module to refine the noisy depth maps, yielding more accurate corrected depth maps. Alongside the corresponding RGB images, camera poses, and segmentation results, they are inputted into the Neural Radiance Fields (NeRF) network [13].
Figure 2: Overview of our proposed method. From the input RGB sequences, depth map, camera pose and segmentation mask can be predicted. Subsequently, the initially estimated depth undergoes correction based on the quadric assumption. Along with the segmentation mask, camera poses, and image frames, the corrected depth are used for optimization of NeRF network. During the 3D reconstruction process, our proposed quadric ray transformer leverages the quadric priors effectively.
@@ 3.2 Quadric depth correction
3.2.1 Quadric surface fitting
We introduce quadric representations for scene components. First, each point is classified by a segmentation network. Using the segmentation mask, we can crop out all the segmented patches in a frame, forming a set of depth patches. Afterward, the depth patches will be projected into 3D spaces with the given calibration matrix. Our proposed quadric fitting is applied to the patches.
To represent the segmented points as quadric surfaces, we define the quadric implicit function as
@@ 3.2.1 Quadric surface fitting
where $\mathbf{x}=\left[x,y,z\right]^{T}$ is the linear term and $\mathbf{q}=\left[x^{2},y^{2},z^{2},xy,yz,xz\right]^{T}$ is the quadric term calculated from $\mathbf{x}$, and $\mathcal{C}_{q}$, $\mathcal{C}_{l}$ and $c$ are the coefficients to be fitted. Following Narunas et al. [35], we define the cost function for the least-square fitting as the sum of squared differences between the quadric equation and the actual $N$ points:
@@
The cost function becomes:
by setting $\nabla\mathbb{C}_{c}=0$ to obtain the optimal $c^{*}$
where $\bar{q}=\frac{1}{N}\sum_{i=1}^{N}q_{i}$ are the quadric term averaged on points in a patch, $\bar{x}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$ is the linear term.
The intermediate variables are defined as follows:
Setting $\nabla\mathbb{C}_{\mathcal{C}_{l}}=0$ gives
$\mathbb{L}~{}\mathcal{C}_{l}^{*}=\mathbb{N}^{T}\mathcal{C}_{q}$ (6)
By substituting $\mathcal{C}_{l}^{*}$ back to Eq. 3, we can obtain
Figure 3: Quadric-based depth correction. Firstly, segmentation masks are predicted from the input RGB image. Subsequently, the initially noisy predicted depth map undergoes re-projection into 3D space, followed by segmentation into patches. For each patch, a quadric surface is fitted, and the depth value is rectified by recalculating the $z$ coordinate through the quadric function.
3.2.2 Depth correction
Through the inverse re-projection using the pixel locations $(u,v)$ and the depth values $d$, a set of 3D points $p$ can be obtained for each segmented patch, where the depth value directly serves as the $z$ coordinate and $K$ is the calibration matrix:
$p=\left[\begin{array}[]{l}X\\ Y\\ Z\end{array}\right]=d\cdot\left(K^{-1}\cdot\left[\begin{array}[]{l}u\\ v\\ 1\end{array}\right]\right)$ (9)
Equipped with the quadric coefficients derived from Section 3.2.1, we possess the capability to rectify inaccuracies present in the initially noisy depth values. By substituting $xy$ back into $f(\mathcal{C},\mathbf{x})$, an equation involving $z$ with other variables as constants is obtained. Solving this equation yields new $z$ values, representing the rectified depth values. This enables the storage of corrected values at their original $(u,v)$ positions, thereby avoiding challenges associated with re-projection errors. This approach enhances the precision and reliability of the 3D representation. The established quadric surface model proves effective in mitigating noise impact on depth information, ultimately improving the fidelity of the reconstructed spatial data.
Figure. 3 illustrates the effectiveness of our proposed quadratically rectified approach. The original patch exhibits drifting points, particularly along the boundary where significant depth value changes occur. However, the proposed approach rectifies these points, fitting them to the quadric surface. A comparison between the corrected and noisy predicted depth maps clearly shows that the boundaries of different patches become more defined, closely resembling the ground truth depth.
@@ 3.3 Quadric Ray transformer
3.3.1 Intra-Ray Transformer
Drawing inspiration from the ray transformer used in IBRNet [38], our approach employs an attention mechanism along the ray to model relationships between sampled points. Initially, rays are constructed from the provided image and calibration matrix, and $N_{s}$ points are uniformly sampled along each ray within the range $[d_{near},d_{far}]$. Utilizing the corrected depth values, denoted as $d$, we refine the sampling process by selecting additional $N_{d}$ samples within the range of $[0.95d,1.05d]$. We do not simply sample around $d$ because potential errors in the corrected depth values might lead to an extended sampling distance away from the true surface.
Following TensoRF [3], we factorize the 4D tensor representing a scene into vectors and matrices. Given the sampled points $p\in R^{B\times(N_{s}+N_{d})\times 3}$, where $B$ is the number of sampled rays, we query the volume density features $f_{\sigma}\in R^{B\times(N_{s}+N_{d})\times D}$ and feed them into the intra-ray transformer. Processing density features instead of color features is a wise choice as they solely depend on $xyz$ positions, simplifying feature aggregation based on the attention mechanism. Additionally, the smaller number of feature channels in density does not significantly increase the computational burden. The density features of each point along the ray are updated using a self-attention module:
@@ 3.3.1 Intra-Ray Transformer
3.3.2 Inter-Ray Transformer
Intuitively, points belonging to the same semantic quadric are more likely to exhibit similar textural and spatial features, while rays sampled from different quadric surfaces might differ from each other in terms of textures and geometries. Therefore, leveraging the updated density features obtained from the intra-ray transformer, which incorporates rich spatial information along the ray, we further propose the inter-ray transformer to capture the relationship across rays.
The inter-ray transformer operates on the updated density features of ray points by incorporating the semantic information. First, the updated density features $f_{\sigma}^{{}^{\prime}}$ are concatenated with the semantic features $f_{s}$. Followed by a fusion network, new density features, $f_{\sigma}^{{}^{\prime\prime}}\in R^{B\times N\times D}$ , which integrate the semantic priors can be obtained.
@@ 3.3.2 Inter-Ray Transformer
As shown in Fig. 4, the inter-ray transformer then takes the transposed density features, $f_{\sigma}^{{}^{\prime\prime}T}\in R^{N\times B\times D}$, as input, where the attention map is calculated across different rays along the $B$ dimension:
@@
where $\delta_{p}^{{}^{\prime}}$ is the positional encoding in attention mechanism. The inter-ray transformer further facilitates feature interaction across rays. Together with the intra-ray transformer, our proposed network achieves feature aggregation across a broad range of ray points, enhancing rendering accuracy by incorporating quadric priors and additional spatial information.

### v2

@@ Abstract
In this paper, we reimagine volumetric representations through the lens of quadrics. We posit that rigid scene components can be effectively decomposed into quadric surfaces. Leveraging this assumption, we reshape the volumetric representations with million of cubes by several quadric planes, which results in more accurate and efficient modeling of 3D scenes in SLAM contexts. First, we use the quadric assumption to rectify noisy depth estimations from RGB inputs. This step significantly improves depth estimation accuracy, and allows us to efficiently sample ray points around quadric planes instead of the entire volume space in previous NeRF-SLAM systems. Second, we introduce a novel quadric-decomposed transformer to aggregate information across quadrics. The quadric semantics are not only explicitly used for depth correction and scene decomposition, but also serve as an implicit supervision signal for the mapping network. Through rigorous experimental evaluation, our method exhibits superior performance over other approaches relying on estimated depth, and achieves comparable accuracy to methods utilizing ground truth depth on both synthetic and real-world datasets.
(a) Volumetric representation
(b) Quadric representation
Figure 1: Comparison between the volume representation and our quadric representation.We partition the 3D scenes into different quadric surfaces. Instead of processing every points independently, we can fully exploit the correlation between points on the same quadric surfaces.
Keywords: Neural Radiance Fields, Simultaneous Localization and Mapping
@@ 1 Introduction
To enable high-quality 3D scene reconstruction, recent SLAM approaches resort to NeRFs and Gaussians as key map representations. Specifically, NeRFs [7], employing neural networks i.e. MLPs, are recognized for their ability to achieve photo-realistic novel-view synthesis, as demonstrated in studies such as iMAP [5], Nice-SLAM [8]. Conversely, 3D Gaussian Splatting (3DGS) [9] relies on dense Gaussian representations to learn scene geometries effectively. Despite achieving photo-realistic novel-view synthesis, NeRFs and Gaussians have a few drawbacks when applied in SLAM systems. On the one hand, MLP-based NeRFs are slow to train while achieving significant memory efficiency. On the other hand, 3D Gaussian Splatting offers near-real-time training capabilities but suffer from large memory consumption and sensitivity to inaccurate camera pose estimation. Additionally, both methods focus on dense high-quality reconstruction, demanding substantial representation efforts to capture fine-grained details.
@@
Addressing these challenges necessitates the design of map representations that are compatible with both learning-based localization front-end and NeRF(GS)-based reconstruction module. Therefore, to enhance scene geometry representation in SLAM, we introduce a novel method, termed Q-SLAM. As illustrated in Figure 1, we integrate the proposed quadric representations across the SLAM pipeline in several key areas.
@@
(1) The well-established front-end tracking module can take RGB stream as input and offer a rough depth estimation while estimating camera poses. Nevertheless, this depth estimation encounters significant challenges, particularly at edges and in texture-less regions, where accuracy tends to degrade. To address these issues, we introduce a quadric- based depth correction module, leveraging quadric surfaces to refine the depth estimation, thus improving overall scene reconstruction accuracy, especially in complex and challenging environments.
@@
(2) In the mapping module, departing from previous NeRF-based SLAM approaches like [4, 8], which rely on volumetric representations, our method transforms densely volumetric scenes - comprising millions of cubes (e.g., $300\times 300\times 300$ voxels) - into a more manageable set of quadric surfaces (50-100 quadric surfaces per keyframe). This approach maintains the integrity of the 3D scene descriptions while significantly reducing complexity, by sampling around the quadric surfaces instead of the whole 3D space.
@@
(3) During the rendering phase, we develop a quadric-ray transformer. We employ importance-sampling based on the quadric-rectified depth values, and we emphasize that sampled points on a quadric naturally belong to the same instance, which facilitates the modeling of their interrelationships during rendering, akin to the method in [10, 11]. We first perform feature interaction between points along a ray to aggregate information and then model the relationships between these surface points across rays using a transformer.
@@
(4) Notably, we extend the utility of quadric semantics beyond mere depth correction and scene decomposition. They are also used as a supervision signal throughout the optimization of the NeRF network to learn robust and accurate features. With the integration of quadrics into the pipeline, we also consider the estimated pose as a learnable parameter in our mapping phase. This introduces a differentiable optimization process to refine the pose estimation, enhancing overall system accuracy.
@@ 2.1 Dense Visual SLAM
Neural Radiance Fields (NeRF) [12] have exhibited notable efficacy in diverse applications, including view synthesis [12, 10, 13] and 3D reconstruction [14, 15]. RGBD SLAM: Innovative contributions such as iMAP [5] and Nice-SLAM [8] have pioneered the integration of NeRF into SLAM systems, demonstrating commendable tracking and mapping performance for indoor scenes. Co-SLAM [1] adopts a multi-resolution hash-grid representation for the scene to expedite convergence. E-SLAM [16] integrates the multi-scale axis-aligned perpendicular feature planes for efficient reconstruction of 3D scenes. By anchoring neural scene features in a point cloud generated iteratively in an input-driven manner, Point-SLAM [17] optimizes runtime and memory usage while maintaining fine detail resolution. Besides using NeRF as the map representation, recent SLAM systems also resort to 3D Gaussian Splatting (3DGS) [9] due to its explicit representation property and real-time rendering performance. SplaTAM [18] pioneers in replacing NeRF with 3DGS and has achieved promising results in both tracking and mapping. GS-SLAM [19] further proposes a coarse-to-fine technique to select reliable 3D Gaussian representations for camera pose optimization effectively. However, these GS-based approaches have one drawbacks in common, much higher requirements for memory usage. All these approaches above rely on ground truth depth map for supervision of NeRF (or 3DGS) training, which is hard to obtain in real-world environment, thereby limiting their applicability in robotics. Conversely, Monocular SLAM method emerges, with only RGB images as input. In particular, NeRF-SLAM [4] integrates DROID-SLAM [20] for pose estimation and depth prediction and use Instant-NGP [21] to fit a NeRF. GO-SLAM [2] globally optimize poses and 3D reconstruction. However, the performance of these methods in 3D reconstruction is impeded by the inherent limitations of noisy predicted depth maps.
@@ 2.2 Scene representations
Volumetric Representations are predominantly utilized in Neural Radiance Fields (NeRFs) [22, 23, 12], and in 3D perception [24, 25, 26]. These representations encapsulate a 3D scene through a volumetric approach, where each volume element embodies occupancy probabilities or task-specific features. However, deriving volumetric representations solely from visual inputs presents significant challenges, particularly in terms of geometric estimation, leading to ambiguous and noisy outputs. Point-Cloud Representation offers a sparser alternative [27]. Point-SLAM [17] utilize a dynamic point density strategy to reduce computational and memory usage. Yet sparse point-clouds struggle to depict scenes comprehensively, while dense point-clouds encounter similar complications to volumetric representations, as previously mentioned. Consequently, few vision-based SLAM systems adopt point-based representations. Additionally, primitives like Planes [28] and Quadric-Meshes [29, 30, 31] have been proposed for scene representation. QuadricSLAM [32] estimates 3D quadric surfaces from 2D multi-view images and uses them as landmark representations for camera tracking in the SLAM pipeline. Dense Planar SLAM [33] uses bounded planes and surfels extracted from depth images to build a real-time SLAM system, but ignore the fine-grained reconstruction of 3D scenes. Point-plane SLAM [34] exploits the constraints of plane edges, which are predominant features that are less affected by measurement noise. ManhattanSLAM [35] utilizes the indoor structural information to estimate camera poses based on the Manhattan World (MW) assumption. However, these are typically employed for regularization during optimization or are challenging to optimize, limiting their applicability in visual SLAM. Drawing inspiration from quadric representations in LiDAR SLAM [3], we introduce quadric representations in visual monocular SLAM.
@@ 3.1 Overview
Our framework, Q-SLAM, as illustrated in Fig. 2, takes monocular RGB sequences as input. Initially, Droid-SLAM [20] predicts rough depth maps and initial camera poses from these inputs. Concurrently, a pretrained segmentation network is employed to estimate segmentation masks from these images. These masks are then utilized by the quadric-based depth correction module to refine the noisy depth maps, yielding more accurate corrected depth maps. Alongside the corresponding RGB images, camera poses, and segmentation results, they are inputted into the NeRF network.
Figure 2: Overview of our proposed method. From the input RGB sequences, we can predict depth map, camera pose and segmentation mask. Subsequently, the initially estimated depth undergoes correction based on the quadric assumption. Along with the segmentation mask, camera poses, and image frames, the corrected depth are used for optimization of NeRF network. During the 3D reconstruction process, our proposed quadric ray transformer leverages the quadric priors effectively.
@@ 3.2 Quadric depth correction
Through the inverse re-projection with the pixel locations $(u,v)$ and the depth values $d$, a set of 3D points $p$ can be obtained for each segmented patch, where the depth value directly serves as the $z$ coordinate and $K$ is the calibration matrix: $[X,Y,Z]^{T}=d\cdot(K^{-1}[u,v,1]^{T})$
To represent the points of each segmented patch as quadric surfaces, we define the quadric implicit function as follows:
@@
where $\mathbf{x}=\left[x,y,z\right]^{T}$ is the linear term and $\mathbf{q}=\left[x^{2},y^{2},z^{2},xy,yz,xz\right]^{T}$ is the quadric term calculated from $\mathbf{x}$, and $\mathcal{C}_{q}$, $\mathcal{C}_{l}$ and $c$ are the coefficients to be fitted. Following Narunas et al. [37], we define the cost function for the least-square fitting as the sum of squared distances between the quadric surface and the actual $N$ points to be fitted:
@@
Minimizing over Eq. 2, we obtain the quadric coefficients. We preserve the patches with fitting error below a predefined threshold, which implies a good fitting surface for the following depth correction. By substituting $(X,Y)$ back into $f(\mathcal{C},\mathbf{x})$, an equation involving $z$ with other variables as constants is obtained. Solving this equation yields new $Z$ values, representing the rectified depth values.
Figure 3: Quadric depth correction
The established quadric surface model proves effective in mitigating noise impact on depth information, ultimately improving the fidelity of the reconstructed spatial data. Fig. 3 illustrates the effectiveness of our proposed quadratically rectified approach. The original patch exhibits drifting points, particularly along the boundary where significant depth value changes occur. However, the proposed approach rectifies these points, fitting them to the quadric surface.
@@ 3.3 Quadric Ray transformer
Intra-Ray Transformer: Drawing inspiration from IBRNet [10], our approach employs an attention mechanism along the ray to model relationships between sampled points. (Fig. 4)
Initially, $N_{s}$ points are uniformly sampled along each ray within the range $[d_{near},d_{far}]$. Utilizing the corrected depth values, denoted as $d$, we refine the sampling process by selecting additional $N_{d}$ samples within the range of $[0.95d,1.05d]$. Given the sampled points $p\in R^{B\times(N_{s}+N_{d})\times 3}$, where $B$ is the number of sampled rays, we query the volume density features $f_{\sigma}\in R^{B\times(N_{s}+N_{d})\times D}$ and feed them into the intra-ray transformer. Processing density features instead of color features is a wise choice as they solely depend on $xyz$ positions, simplifying feature aggregation based on the attention mechanism. Additionally, the smaller number of feature channels in density does not significantly increase the computational burden, resulting in an efficient representation for the application of SLAM.
The density features of each point along the ray are updated using self-attention:
@@
Inter-Ray Transformer: Intuitively, points belonging to the same semantic quadric are more likely to exhibit similar textural and spatial features, while rays sampled from different quadric surfaces might differ from each other in terms of textures and geometries. Therefore, we further propose the inter-ray transformer to capture the relationship across rays.
The inter-ray transformer operates on the updated density features from intra-ray transformer. First, the updated density features $f_{\sigma}^{{}^{\prime}}$ are concatenated with the semantic features $f_{s}$. Followed by a fusion network, new density features with semantic priors, $f_{\sigma}^{{}^{\prime\prime}}\in R^{B\times N\times D}$ , can be obtained.
@@
As shown in Fig. 4, the inter-ray transformer then takes the transposed density features, $f_{\sigma}^{{}^{\prime\prime}T}\in R^{N\times B\times D}$, as input, where attention map is calculated across different rays along the $B$ dimension:
@@
where $\delta_{p}^{{}^{\prime}}$ is the positional encoding in attention mechanism.
The inter-ray transformer further facilitates feature interaction across rays. This process allows surface points to capture more inter-ray information from other rays from the same quadric surface. Together with the intra-ray transformer, our proposed network achieves feature aggregation across a broad range of ray points, enhancing rendering accuracy by incorporating quadric priors and additional spatial information.